OpenStudy (anonymous):

Help me? |x + 3| = 3

OpenStudy (anonymous):

A. {0, 6} B. {0, -6} C. {-6, 6} D. {-6} These are the choices. How are you getting that?

OpenStudy (anonymous):

Buttt hoooowww

OpenStudy (skullpatrol):

What does absolute value mean?

OpenStudy (pinknabastak):

and -6 +3 = -3 and the absoulute value of -3 is 3

OpenStudy (anonymous):

There can be more answers not just 0.

OpenStudy (skullpatrol):

@mconnors please give me the definition of absolute value?

OpenStudy (bgrg007):

Cuz to find x, you gotta keep it separate from the other regular numbers. So you take away 3(subtract 3) on both sides, which gets u x=0. Also, another form of absolute value of 3 is -3. so u know its -6 that'll give u -3 at the end.

OpenStudy (bgrg007):

i actually got reported with a warning for giving u the right answer. wtf lmaoo

OpenStudy (anonymous):

thank you! @skullpatrol : the magnitude of a real number without regard to its sign.

OpenStudy (skullpatrol):

OK what does the equation say?

OpenStudy (pinknabastak):

he already got the answer...

OpenStudy (skullpatrol):

|x + 3| = 3

OpenStudy (skullpatrol):

@mconnors whatever number is on the left hand side must be the same as the the number on the right hand side, correct?

OpenStudy (skullpatrol):

Otherwise it is not a true statement.

OpenStudy (skullpatrol):

"=" means have the same value @mconnors

OpenStudy (skullpatrol):

@mconnors are you still there?

OpenStudy (shamil98):

|x + 3| = 3 means x + 3 = 3 -(x+3) = 3 OR simply: x+3 = -3 from there just solve normally.

OpenStudy (nincompoop):

in a number line from a particular point, x, when three units are added we get three. since x could either be a positive or a negative, we generally count the units from the right side towards x and left side towards x. 

OpenStudy (nincompoop):



OpenStudy (nincompoop):

upon understanding the visual representation, follow shamil's explanation to solve algebraically.